OpenStudy (anonymous):

HELP ! What is the sum of the first n terms of this series? 4 + 6 + 8 + 10 + 12 + ⋯

OpenStudy (anonymous):

answer is 8

OpenStudy (shamim):

common difference d=?

OpenStudy (anonymous):

no its not @jot

OpenStudy (anonymous):

why

OpenStudy (anonymous):

\[\Sigma \ its and answer whith this

OpenStudy (anonymous):

\[\Sigma \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@iGreen @Kemist @ParthKohli helppp

OpenStudy (anonymous):

see??

OpenStudy (anonymous):

@iGreen do u know this?

OpenStudy (igreen):

No, sorry..

OpenStudy (anonymous):

;'O

OpenStudy (anonymous):

@ParthKohli do you know?

OpenStudy (tkhunny):

4 + 6 + 8 + 10 + 12 + ⋯ 2*(2 + 3 + 4 + 5 + 6 + ⋯) The odd fact you should memorize is this, 1 + 2 + 3 + ... + n = \(\dfrac{n(n+1)}{2}\) Check out inside those parentheses. Do you see a relationship?

OpenStudy (anonymous):

not really 0.0 :'0

OpenStudy (tkhunny):

1 + 2 + 3 + 4 + ... + n -- n terms starting with 1. Sum is \(\dfrac{n(n+1)}{2}\) 2 + 3 + 4 + ... + n + (n+1) -- n terms starting with 2. Sum is \(\dfrac{(n+1)(n+2)}{2} - 1\) Final answer is what? 4 + 6 + 8 + 10 + 12 + ⋯ 2*(2 + 3 + 4 + 5 + 6 + ⋯)

OpenStudy (anonymous):

@geerky42 ??

OpenStudy (anonymous):

@tkhunny I DONT KNOW D:<

OpenStudy (tkhunny):

\(2\left(\dfrac{(n+1)(n+2)}{2} + 1\right) = (n+1)(n+2) + 2\)

OpenStudy (anonymous):